A man earns a salary plus 10% a year bonus. If his bonus one year was $1,795, what was his total pay for the year?
Answer by dkppathak(439)

let his salary is x rupees
10 percent of salary is vonus =x/10
x/10=1795
x=17950
total pay= salary + vonus =17950+1795=19745
toala pay RS 17945